OpenStudy (anonymous):

Two lines, R and M, are represented by the following equations: Line R: 2x + 2y = 18 Line M: x + y = 9 Which statement is true about the solution to the set of equations? It is (18, 9). It is (9, 18). There are infinitely many solutions. There is no solution.

OpenStudy (anonymous):

I still need help with this!

OpenStudy (anonymous):

what do u think the answer is ?

OpenStudy (anonymous):

hint solve simultaneously

OpenStudy (anonymous):

D. there is no solution if not then A. (18, 9)

OpenStudy (anonymous):

yes u are right its D because

OpenStudy (anonymous):

2x + 2y = 18 x + y = 9 solving simultaneously 1(2x+2y=18) 2(x+y=9) 2x+2y=18 2x+2y=18 you cant eliminate one of those letters, and while keeping the other letter (either x or y)

OpenStudy (anonymous):

I have two more questions I need help with, if you are willing to help out.

OpenStudy (anonymous):

yea sure plz post em on new thread :)

OpenStudy (anonymous):

Gotcha